\begin{thebibliography}{18}%
\makeatletter
\providecommand \@ifxundefined [1]{%
 \@ifx{#1\undefined}
}%
\providecommand \@ifnum [1]{%
 \ifnum #1\expandafter \@firstoftwo
 \else \expandafter \@secondoftwo
 \fi
}%
\providecommand \@ifx [1]{%
 \ifx #1\expandafter \@firstoftwo
 \else \expandafter \@secondoftwo
 \fi
}%
\providecommand \natexlab [1]{#1}%
\providecommand \enquote  [1]{``#1''}%
\providecommand \bibnamefont  [1]{#1}%
\providecommand \bibfnamefont [1]{#1}%
\providecommand \citenamefont [1]{#1}%
\providecommand \href@noop [0]{\@secondoftwo}%
\providecommand \href [0]{\begingroup \@sanitize@url \@href}%
\providecommand \@href[1]{\@@startlink{#1}\@@href}%
\providecommand \@@href[1]{\endgroup#1\@@endlink}%
\providecommand \@sanitize@url [0]{\catcode `\\12\catcode `\$12\catcode
  `\&12\catcode `\#12\catcode `\^12\catcode `\_12\catcode `\%12\relax}%
\providecommand \@@startlink[1]{}%
\providecommand \@@endlink[0]{}%
\providecommand \url  [0]{\begingroup\@sanitize@url \@url }%
\providecommand \@url [1]{\endgroup\@href {#1}{\urlprefix }}%
\providecommand \urlprefix  [0]{URL }%
\providecommand \Eprint [0]{\href }%
\providecommand \doibase [0]{https://doi.org/}%
\providecommand \selectlanguage [0]{\@gobble}%
\providecommand \bibinfo  [0]{\@secondoftwo}%
\providecommand \bibfield  [0]{\@secondoftwo}%
\providecommand \translation [1]{[#1]}%
\providecommand \BibitemOpen [0]{}%
\providecommand \bibitemStop [0]{}%
\providecommand \bibitemNoStop [0]{.\EOS\space}%
\providecommand \EOS [0]{\spacefactor3000\relax}%
\providecommand \BibitemShut  [1]{\csname bibitem#1\endcsname}%
\let\auto@bib@innerbib\@empty
%</preamble>
\bibitem [{\citenamefont {Henkes}\ and\ \citenamefont
  {Chakraborty}(2009)}]{Bulbul_pre_2009}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {S.}~\bibnamefont
  {Henkes}}\ and\ \bibinfo {author} {\bibfnamefont {B.}~\bibnamefont
  {Chakraborty}},\ }\bibfield  {title} {\bibinfo {title} {Statistical mechanics
  framework for static granular matter},\ }\href
  {https://doi.org/10.1103/PhysRevE.79.061301} {\bibfield  {journal} {\bibinfo
  {journal} {Phys. Rev. E}\ }\textbf {\bibinfo {volume} {79}},\ \bibinfo
  {pages} {061301} (\bibinfo {year} {2009})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Chowdhury}\ \emph {et~al.}(2016)\citenamefont
  {Chowdhury}, \citenamefont {Abraham}, \citenamefont {Hudson},\ and\
  \citenamefont {Harrowell}}]{Harrowell_jcp_stress_correlations_2016}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {S.}~\bibnamefont
  {Chowdhury}}, \bibinfo {author} {\bibfnamefont {S.}~\bibnamefont {Abraham}},
  \bibinfo {author} {\bibfnamefont {T.}~\bibnamefont {Hudson}},\ and\ \bibinfo
  {author} {\bibfnamefont {P.}~\bibnamefont {Harrowell}},\ }\bibfield  {title}
  {\bibinfo {title} {Long range stress correlations in the inherent structures
  of liquids at rest},\ }\href {https://doi.org/10.1063/1.4944620} {\bibfield
  {journal} {\bibinfo  {journal} {J. Chem. Phys.}\ }\textbf {\bibinfo {volume}
  {144}},\ \bibinfo {pages} {124508} (\bibinfo {year} {2016})}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont
  {Lema\^{\i}tre}(2017)}]{lemaitre_pre_2017_stress_correlations_2D}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {A.}~\bibnamefont
  {Lema\^{\i}tre}},\ }\bibfield  {title} {\bibinfo {title} {Inherent stress
  correlations in a quiescent two-dimensional liquid: Static analysis including
  finite-size effects},\ }\href {https://doi.org/10.1103/PhysRevE.96.052101}
  {\bibfield  {journal} {\bibinfo  {journal} {Phys. Rev. E}\ }\textbf {\bibinfo
  {volume} {96}},\ \bibinfo {pages} {052101} (\bibinfo {year}
  {2017})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont
  {Lema\^{\i}tre}(2018)}]{lemaitre_jcp_2018_stress_correlations_3D}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {A.}~\bibnamefont
  {Lema\^{\i}tre}},\ }\bibfield  {title} {\bibinfo {title} {Stress correlations
  in glasses},\ }\href {https://doi.org/10.1063/1.5041461} {\bibfield
  {journal} {\bibinfo  {journal} {J. Chem. Phys.}\ }\textbf {\bibinfo {volume}
  {149}},\ \bibinfo {pages} {104107} (\bibinfo {year} {2018})}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont
  {DeGiuli}(2018{\natexlab{a}})}]{eric_d_field_theory_prl_2018}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {E.}~\bibnamefont
  {DeGiuli}},\ }\bibfield  {title} {\bibinfo {title} {Field theory for
  amorphous solids},\ }\href {https://doi.org/10.1103/PhysRevLett.121.118001}
  {\bibfield  {journal} {\bibinfo  {journal} {Phys. Rev. Lett.}\ }\textbf
  {\bibinfo {volume} {121}},\ \bibinfo {pages} {118001} (\bibinfo {year}
  {2018}{\natexlab{a}})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont
  {DeGiuli}(2018{\natexlab{b}})}]{eric_d_field_theory_pre_2018}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {E.}~\bibnamefont
  {DeGiuli}},\ }\bibfield  {title} {\bibinfo {title} {Edwards field theory for
  glasses and granular matter},\ }\href
  {https://doi.org/10.1103/PhysRevE.98.033001} {\bibfield  {journal} {\bibinfo
  {journal} {Phys. Rev. E}\ }\textbf {\bibinfo {volume} {98}},\ \bibinfo
  {pages} {033001} (\bibinfo {year} {2018}{\natexlab{b}})}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Shimada}\ and\ \citenamefont
  {De~Giuli}(2020)}]{eric_Shimada_field_theory_arXiv_2020}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {M.}~\bibnamefont
  {Shimada}}\ and\ \bibinfo {author} {\bibfnamefont {E.}~\bibnamefont
  {De~Giuli}},\ }\bibfield  {title} {\bibinfo {title} {Random quench predicts
  universal properties of amorphous solids},\ }\href
  {https://arxiv.org/abs/2008.11896} {\bibfield  {journal} {\bibinfo  {journal}
  {arXiv preprint arXiv:2008.11896}\ } (\bibinfo {year} {2020})}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Tong}\ \emph {et~al.}(2020)\citenamefont {Tong},
  \citenamefont {Sengupta},\ and\ \citenamefont
  {Tanaka}}]{tanaka_emergent_solidity_nat_comm_2020}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {H.}~\bibnamefont
  {Tong}}, \bibinfo {author} {\bibfnamefont {S.}~\bibnamefont {Sengupta}},\
  and\ \bibinfo {author} {\bibfnamefont {H.}~\bibnamefont {Tanaka}},\
  }\bibfield  {title} {\bibinfo {title} {Emergent solidity of amorphous
  materials as a consequence of mechanical self-organisation},\ }\href
  {https://doi.org/10.1038/s41467-020-18663-7} {\bibfield  {journal} {\bibinfo
  {journal} {Nat. Commun.}\ }\textbf {\bibinfo {volume} {11}},\ \bibinfo
  {pages} {4863} (\bibinfo {year} {2020})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Nampoothiri}\ \emph {et~al.}(2020)\citenamefont
  {Nampoothiri}, \citenamefont {Wang}, \citenamefont {Ramola}, \citenamefont
  {Zhang}, \citenamefont {Bhattacharjee},\ and\ \citenamefont
  {Chakraborty}}]{Bulbul_prl_2020_emergent_elasticity}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {J.~N.}\ \bibnamefont
  {Nampoothiri}}, \bibinfo {author} {\bibfnamefont {Y.}~\bibnamefont {Wang}},
  \bibinfo {author} {\bibfnamefont {K.}~\bibnamefont {Ramola}}, \bibinfo
  {author} {\bibfnamefont {J.}~\bibnamefont {Zhang}}, \bibinfo {author}
  {\bibfnamefont {S.}~\bibnamefont {Bhattacharjee}},\ and\ \bibinfo {author}
  {\bibfnamefont {B.}~\bibnamefont {Chakraborty}},\ }\bibfield  {title}
  {\bibinfo {title} {Emergent elasticity in amorphous solids},\ }\href
  {https://doi.org/10.1103/PhysRevLett.125.118002} {\bibfield  {journal}
  {\bibinfo  {journal} {Phys. Rev. Lett.}\ }\textbf {\bibinfo {volume} {125}},\
  \bibinfo {pages} {118002} (\bibinfo {year} {2020})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Calladine}(1978)}]{calladine_1978}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {C.}~\bibnamefont
  {Calladine}},\ }\bibfield  {title} {\bibinfo {title} {Buckminster fullers
  tensegrity structures and clerk maxwell's rules for the construction of stiff
  frames},\ }\href
  {https://doi.org/https://doi.org/10.1016/0020-7683(78)90052-5} {\bibfield
  {journal} {\bibinfo  {journal} {Int. J. Solids Struct.}\ }\textbf {\bibinfo
  {volume} {14}},\ \bibinfo {pages} {161 } (\bibinfo {year}
  {1978})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Lerner}\ \emph {et~al.}(2012)\citenamefont {Lerner},
  \citenamefont {D{\"u}ring},\ and\ \citenamefont {Wyart}}]{asm_pnas_2012}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {E.}~\bibnamefont
  {Lerner}}, \bibinfo {author} {\bibfnamefont {G.}~\bibnamefont {D{\"u}ring}},\
  and\ \bibinfo {author} {\bibfnamefont {M.}~\bibnamefont {Wyart}},\ }\bibfield
   {title} {\bibinfo {title} {A unified framework for non-brownian suspension
  flows and soft amorphous solids},\ }\href
  {https://doi.org/10.1073/pnas.1120215109} {\bibfield  {journal} {\bibinfo
  {journal} {Proc. Natl. Acad. Sci. U.S.A.}\ }\textbf {\bibinfo {volume}
  {109}},\ \bibinfo {pages} {4798} (\bibinfo {year} {2012})}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Lerner}(2018)}]{states_of_self_stress_epje_2018}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {E.}~\bibnamefont
  {Lerner}},\ }\bibfield  {title} {\bibinfo {title} {Quasilocalized states of
  self stress in packing-derived networks},\ }\href
  {https://doi.org/10.1140/epje/i2018-11705-9} {\bibfield  {journal} {\bibinfo
  {journal} {Eur. Phys. J. E}\ }\textbf {\bibinfo {volume} {41}},\ \bibinfo
  {pages} {93} (\bibinfo {year} {2018})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Snoeijer}\ \emph {et~al.}(2004)\citenamefont
  {Snoeijer}, \citenamefont {Vlugt}, \citenamefont {van Hecke},\ and\
  \citenamefont {van Saarloos}}]{vanSaarloos_prl_2004}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {J.~H.}\ \bibnamefont
  {Snoeijer}}, \bibinfo {author} {\bibfnamefont {T.~J.~H.}\ \bibnamefont
  {Vlugt}}, \bibinfo {author} {\bibfnamefont {M.}~\bibnamefont {van Hecke}},\
  and\ \bibinfo {author} {\bibfnamefont {W.}~\bibnamefont {van Saarloos}},\
  }\bibfield  {title} {\bibinfo {title} {Force network ensemble: A new approach
  to static granular matter},\ }\href
  {https://doi.org/10.1103/PhysRevLett.92.054302} {\bibfield  {journal}
  {\bibinfo  {journal} {Phys. Rev. Lett.}\ }\textbf {\bibinfo {volume} {92}},\
  \bibinfo {pages} {054302} (\bibinfo {year} {2004})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Sussman}\ \emph {et~al.}(2016)\citenamefont
  {Sussman}, \citenamefont {Goodrich},\ and\ \citenamefont
  {Liu}}]{sussman_soft_matter_2016}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {D.~M.}\ \bibnamefont
  {Sussman}}, \bibinfo {author} {\bibfnamefont {C.~P.}\ \bibnamefont
  {Goodrich}},\ and\ \bibinfo {author} {\bibfnamefont {A.~J.}\ \bibnamefont
  {Liu}},\ }\bibfield  {title} {\bibinfo {title} {Spatial structure of states
  of self stress in jammed systems},\ }\href
  {https://doi.org/10.1039/C6SM00094K} {\bibfield  {journal} {\bibinfo
  {journal} {Soft Matter}\ }\textbf {\bibinfo {volume} {12}},\ \bibinfo {pages}
  {3982} (\bibinfo {year} {2016})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Wyart}(2005)}]{matthieu_thesis}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {M.}~\bibnamefont
  {Wyart}},\ }\bibfield  {title} {\bibinfo {title} {On the rigidity of
  amorphous solids},\ }\href {https://doi.org/10.1051/anphys:2006003}
  {\bibfield  {journal} {\bibinfo  {journal} {Ann. Phys. Fr.}\ }\textbf
  {\bibinfo {volume} {30}},\ \bibinfo {pages} {1} (\bibinfo {year}
  {2005})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Lerner}(2017)}]{comment_sss_soft_matter_2017}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {E.}~\bibnamefont
  {Lerner}},\ }\bibfield  {title} {\bibinfo {title} {Comment on ``spatial
  structure of states of self stress in jammed systems'' by d. m. sussman{,} c.
  p. goodrich{,} and a. j. liu{,} soft matter{,} 2016{,} 12{,} 3982},\ }\href
  {https://doi.org/10.1039/C6SM01111J} {\bibfield  {journal} {\bibinfo
  {journal} {Soft Matter}\ }\textbf {\bibinfo {volume} {13}},\ \bibinfo {pages}
  {1530} (\bibinfo {year} {2017})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Lerner}\ \emph {et~al.}(2014)\citenamefont {Lerner},
  \citenamefont {DeGiuli}, \citenamefont {During},\ and\ \citenamefont
  {Wyart}}]{breakdown}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {E.}~\bibnamefont
  {Lerner}}, \bibinfo {author} {\bibfnamefont {E.}~\bibnamefont {DeGiuli}},
  \bibinfo {author} {\bibfnamefont {G.}~\bibnamefont {During}},\ and\ \bibinfo
  {author} {\bibfnamefont {M.}~\bibnamefont {Wyart}},\ }\bibfield  {title}
  {\bibinfo {title} {Breakdown of continuum elasticity in amorphous solids},\
  }\href {https://doi.org/10.1039/C4SM00311J} {\bibfield  {journal} {\bibinfo
  {journal} {Soft Matter}\ }\textbf {\bibinfo {volume} {10}},\ \bibinfo {pages}
  {5085} (\bibinfo {year} {2014})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Gartner}\ and\ \citenamefont
  {Lerner}(2016)}]{SciPost2016}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {L.}~\bibnamefont
  {Gartner}}\ and\ \bibinfo {author} {\bibfnamefont {E.}~\bibnamefont
  {Lerner}},\ }\bibfield  {title} {\bibinfo {title} {{Nonlinear modes
  disentangle glassy and Goldstone modes in structural glasses}},\ }\href
  {https://doi.org/10.21468/SciPostPhys.1.2.016} {\bibfield  {journal}
  {\bibinfo  {journal} {SciPost Phys.}\ }\textbf {\bibinfo {volume} {1}},\
  \bibinfo {pages} {016} (\bibinfo {year} {2016})}\BibitemShut {NoStop}%
\end{thebibliography}

\documentclass[aps,twocolumn,pre,superscriptaddress, floatfix]{revtex4-2}

\usepackage{xcolor,hyperref}
\hypersetup{
   colorlinks,
   linkcolor={blue!50!black},
   citecolor={blue!50!black},
   urlcolor={blue!80!black}
}

\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{color}

\usepackage{graphicx}

\DeclareMathAlphabet{\mathitbf}{OML}{cmm}{b}{it}

\DeclareMathAlphabet{\mathsfit}{T1}{\sfdefault}{\mddefault}{\sldefault}
\SetMathAlphabet{\mathsfit}{bold}{T1}{\sfdefault}{\bfdefault}{\sldefault}

\newcommand{\zerovector}{\mathBold 0}
\newcommand{\ket}[1]{|#1\rangle}

\newcommand{\xv}{\mathitbf x}

\newcommand{\calBold}[1]{\mbox{\boldmath${\cal #1}$}}
\newcommand{\mathBold}[1]{\mbox{\boldmath$#1$}}
\newcommand{\dbar}{{\,\mathchar'26\mkern-12mu d}}

\begin{document}

\title{\fontsize{11}{3} \selectfont Note: Simple argument for emergent anisotropic stress correlations in disordered solids}

\author{Edan Lerner}
\email{e.lerner@uva.nl}
\affiliation{Institute for Theoretical Physics, University of Amsterdam, Science Park 904, Amsterdam, Netherlands}

\maketitle

It is now well-established that mechanical equilibrium in athermal disordered solids gives rise to anisotropic spatial correlations of the coarse-grained stress field \cite{Bulbul_pre_2009,Harrowell_jcp_stress_correlations_2016,lemaitre_pre_2017_stress_correlations_2D,lemaitre_jcp_2018_stress_correlations_3D,eric_d_field_theory_prl_2018,eric_d_field_theory_pre_2018,eric_Shimada_field_theory_arXiv_2020,tanaka_emergent_solidity_nat_comm_2020,Bulbul_prl_2020_emergent_elasticity} that decay in space as $1/r^{\dbar}$, where $r$ is the distance from the origin, and $\dbar$ denotes the spatial dimension. In this note we present a simple, geometry based argument for the scaling form of the emergent spatial correlations of the stress field in disordered solids. The presented approach bears some conceptual similarities with the field-theoretic approach of \cite{eric_d_field_theory_prl_2018,eric_d_field_theory_pre_2018,eric_Shimada_field_theory_arXiv_2020}. 

Consider a disordered solid whose constituent particles interact via pairwise, radially-symmetric interactions. We denote particle coordinates by $\xv_i$, the radius vector between a pair $i,j$ of interacting particles by $\xv_{ij}\!\equiv\!\xv_j\!-\!\xv_i$, and the distance between them by $r_{ij}\!\equiv\!\sqrt{\xv_{ij}\!\cdot\!\xv_{ij}}$. With these definitions, the magnitude of the forces $f_{ij}$ between particle $i$ and particle $j$ are defined to be positive (negative) if they are repulsive (attractive), and the condition of mechanical equilibrium reads
\begin{equation}\label{foo00}
\calBold{F}_i = \sum_j f_{ij} \xv_{ij}/r_{ij} = \zerovector\,,
\end{equation}
where $\calBold{F}_i$ denotes the \emph{net} force on the $i$'th particle.

Eq.~\eqref{foo00} can be cast into a more convenient, bra-ket notation in the form \cite{calladine_1978,asm_pnas_2012,states_of_self_stress_epje_2018}
\begin{equation}
\ket{\calBold{F}} = {\cal S}^T\ket{f} = \zerovector\,,
\end{equation}
where we have defined the linear operator ${\cal S}$ with components
\begin{equation}\label{foo04}
{\cal S}_{ij,k} \equiv \frac{\partial r_{ij}}{\partial\xv_k} = (\delta_{jk} - \delta_{ik})\frac{\xv_{ij}}{r_{ij}}\,,
\end{equation}
which takes vectors from the space of particles' coordinates, to the space of pairs of interacting particles ($\delta_{jk}$ is the Kronecker delta). ${\cal S}$ is constructed based on \emph{geometric} information of the network of pairwise interactions of a disordered configuration. States occupying the null-space of ${\cal S}^T$ --- which result in zero net force $\calBold{F}_i\!=\!\zerovector$ on each particle --- are known as \emph{states of self stress} \cite{vanSaarloos_prl_2004,sussman_soft_matter_2016,states_of_self_stress_epje_2018}. The number of orthogonal states of self stress of a given configuration scales as $z\!-\!2\dbar$ where $z$ is the mean coordination of the network of interactions \cite{sussman_soft_matter_2016}.

Consider next a disordered solid configuration, and assign any set of \emph{random} forces $\phi_\alpha$ between the particles, which obey the statistics
\begin{equation}\label{foo01}
\overline{\phi_\alpha} = 0\,, \quad\mbox{and}\quad \overline{\phi_\alpha\phi_\beta} = \delta_{\alpha\beta}\,,
\end{equation}
where here and in what follows Greek letters label \emph{pairs} of interacting particles, $\overline{\bullet}$ denotes a spatial average, $\delta_{\alpha\beta}\!=\!1$ if the pair $\alpha$ is equal to the pair $\beta$, and $\delta_{\alpha\beta}\!=\!0$ otherwise. The assigned `ancestral' forces $\phi_\alpha$ are supposed to represent the random forces between particles of a liquid before vitrification, shown recently \cite{tanaka_emergent_solidity_nat_comm_2020} to be uncorrelated in liquids residing just above their respective glass transition temperatures. Since the forces $\phi_\alpha$ are random, then clearly ${\cal S}^T\ket{\phi}\!\ne\!\zerovector$, i.e.~the system is out of mechanical equilibrium. We note that the $\delta$-correlated ancestral forces requirement is analogous to the integrability condition on the pressure autocorrelation as discussed in Refs.~\cite{Bulbul_pre_2009,lemaitre_pre_2017_stress_correlations_2D,lemaitre_jcp_2018_stress_correlations_3D,eric_d_field_theory_prl_2018,eric_d_field_theory_pre_2018,eric_Shimada_field_theory_arXiv_2020}.

We now project the set of random forces $\ket{\phi}$ onto the null space of ${\cal S}^T$, to obtain a state of self stress $\ket{f}$, namely
\begin{equation}
\ket{f} = {\cal P}\ket{\phi}\,, \quad\mbox{such that}\quad {\cal S}^T\ket{f}=\zerovector\,.
\end{equation}
As mentioned above regarding ${\cal S}$, the projection operator ${\cal P}$ is also constructed using geometric information of the network of interactions alone; it is given by \cite{matthieu_thesis,states_of_self_stress_epje_2018,comment_sss_soft_matter_2017}
\begin{equation}\label{foo03}
{\cal P} \equiv {\cal I} - {\cal S}\big({\cal S}^T{\cal S}\big)^{-1}{\cal S}^T\,,
\end{equation}
where ${\cal I}_{\alpha\beta}\!\equiv\!\delta_{\alpha\beta}$ is the interacting-pairs-space identity operator. It is easy to show that ${\cal P}\!=\!{\cal P}^T\! =\! {\cal P}^2$. ${\cal P}$ is, in essence, the discrete analog of the ``transverse projector" operator described in \cite{eric_d_field_theory_prl_2018,eric_d_field_theory_pre_2018}. 

Finally, consider the spatial force correlations
\begin{equation}
\overline{f_\alpha f_\beta} = \overline{\phi_{\kappa}{\cal P}_{\kappa\alpha}^T{\cal P}_{\beta\chi}\phi_{\chi}}\,,
\end{equation}
where repeated indices are understood as summed over. Since the random ancestral forces $\phi_\alpha$ are uncorrelated with the interaction network's geometry encoded in ${\cal P}$, then $\overline{\phi_{\kappa}{\cal P}_{\kappa\alpha}^T{\cal P}_{\beta\chi}\phi_{\chi}}=\overline{{\cal P}_{\kappa\alpha}^T{\cal P}_{\beta\chi}}\,\overline{\phi_{\chi}\phi_{\kappa}}$. Using Eq.~(\ref{foo01}), we find
\begin{equation}\label{foo02}
\overline{f_\alpha f_\beta} = \overline{{\cal P}_{\kappa\alpha}^T{\cal P}_{\beta\chi}}\delta_{\kappa\chi} = \overline{{\cal P}_{\alpha\beta}}\,.
\end{equation}
We point out that (i) the elastic Green's function of a relaxed Hookean spring network derived from any solid's network of pairwise interactions is given by $({\cal S}^T{\cal S})^{-1}$ (assuming unit stiffnesses and masses), as shown in \cite{asm_pnas_2012,states_of_self_stress_epje_2018}. (ii)~${\cal P}_{\alpha\beta}$ (with $\alpha\!\ne\!\beta$) is precisely the displacement response --- projected onto the radius vector of the $\beta$ pair (see definition above) --- that results from applying a unit force dipole on the pair $\alpha$ of the corresponding mapped elastic Hookean spring network \cite{states_of_self_stress_epje_2018}.

According to continuum elasticity theory, the displacement response to a force dipole decays as $1/r^{\dbar-1}$ at large distances $r$ from the force dipole. The average projection of such a displacement response onto the radius vector of an interacting pair located away from the applied force dipole is akin to taking the displacement-response's spatial gradient (due to the difference $\delta_{jk}\!-\!\delta_{ik}$ as appears in Eq.~(\ref{foo04}), see also~\cite{breakdown, SciPost2016}). Therefore $\overline{f_\alpha f_\beta}\!\sim\!\overline{{\cal P}_{\alpha\beta}}\!\sim\! 1/r^{\dbar}$, in agreement with previous work regarding stress correlations in disordered solids~\cite{Bulbul_pre_2009,Harrowell_jcp_stress_correlations_2016,lemaitre_pre_2017_stress_correlations_2D,lemaitre_jcp_2018_stress_correlations_3D,eric_d_field_theory_prl_2018,eric_d_field_theory_pre_2018,tanaka_emergent_solidity_nat_comm_2020,Bulbul_prl_2020_emergent_elasticity,eric_Shimada_field_theory_arXiv_2020}. We note that the anisotropy seen in far-field stress correlations ($\sim\!\cos(4\theta)$ in two-dimensions, with $\theta$ the azimuthal angle) can also be obtained via the aformentioned analogy to the elastic Green's function and the continuum-elastic response to local force dipoles. 

Interestingly, the same exercise can be repeated assuming random ancestral forces $\phi_\alpha$ in a periodic lattice of rigid struts. In that case, one finds $\overline{f_\alpha f_\beta}\!\sim\! {\cal P}_{\alpha\beta}\!\sim\! r^{-\dbar}$ (no spatial average over ${\cal P}$), where ${\cal P}$ is exactly obtained via Eq.~(\ref{foo03}), given any lattice geometry. This means that \emph{structural} disorder itself --- and therefore also isotropy --- are not necessary ingredients for the form of emergent spatial correlations of coarse-grained stress fields.

We thank Gustavo D\"uring, Eran Bouchbinder and Eric DeGiuli for discussions. Financial support from the NWO (Vidi grant no.~680-47-554/3259) is acknowledged. 

\section*{Data availability}
Data sharing is not applicable to this article as no new data were created or analyzed in this study.

\end{document}